OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid $130. This year, he paid $169. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (stormfire1):

(1 - (130/169) ) * 100 = 23%

OpenStudy (anonymous):

Can you explain to me how you got 1 at the beginning?

OpenStudy (stormfire1):

Crap...I made a dumb mistake. Here: 130x = 169 so it should have been (169/130) - 100 = 30 % :)

OpenStudy (stormfire1):

Just in case it's not clear, you have to subtract 100 because 169 is 130% of 130.

OpenStudy (anonymous):

\[\left(\frac{169-130}{130}*100\right)\%=30\% \]